OpenStudy (anonymous):

4/5 of the people at the circus are kids. If 575 tickets were sold, how many are expected to be kids? (460, 115, 1,035)

OpenStudy (anonymous):

4/5 of 575 = 4 * 575 ------ 5 = 4 * 115 = 460